3 square root x -5=3
What does x= ?
3 (square root x) -5 = 3
3 (square root x) = 8
square root x = 8/3
x 64/9

Found 2 solutions by Alan3354, gonzo:
Answer by Alan3354(21608)

3 square root x -5=3
-----------
Is it 3*sqrt(x-5) = 3?
---------------
square both sides
9*(x-5) = 9
x-5 = 1
x = 6

Answer by gonzo(654)